# Universal 3-Edge-Weightability of Regular Graphs

### v1

@@ Abstract.
A graph $G=(V,E)$ is universally $3$-edge-weightable if for every set $\{a,b,c\}$ of three distinct real numbers, there exists an edge-weighting $w:E\to\{a,b,c\}$ such that $d_{w}(u)\neq d_{w}(v)$ for all $uv\in E$. The 1-2-3 Conjecture, recently settled by Keusch, ensures that all nice graphs admit a proper $\{1,2,3\}$-edge-weighting. For regular graphs, this implies proper edge-weightings for arithmetic progressions $\{a,b,c\}$ via affine transformation, but not for general triples due to the spacing-ratio preservation property of affine maps. We prove that every $k$-regular graph with $k\geq 2$ is universally $3$-edge-weightable. This establishes proper edge-weightings for all triples $\{a,b,c\}$, not just arithmetic progressions, showing that regular graphs satisfy a strictly stronger weighting property than what follows from the 1-2-3 Conjecture alone. The proof adapts the independent-set decomposition framework to edge-weightings and introduces a two-phase conflict-resolution scheme that operates independently of the specific choice of $\{a,b,c\}$.
@@ 1. Introduction
All graphs considered are simple, finite, and undirected. A graph is nice if none of its components is isomorphic to $K_{2}$. For a graph $G=(V,E)$ and a subset $Q\subset\mathbb{R}$, a mapping $w:E\to Q$ is called a $Q$-edge weighting of $G$; the weighted degree of a vertex $v$ is defined as $d_{w}(v)=\sum_{e\ni v}w(e)$; an edge $e=uv$ is a conflict if $d_{w}(u)=d_{w}(v)$. If there is no conflict under $w$, then $w$ is called proper. We say $G$ is $Q$-edge-weightable if it admits a proper $Q$-edge weighting, and call $G$ universally $k$-edge-weightable if it is $Q$-edge weightable for every $k$-element subset $Q\subset\mathbb{R}$.
The 1-2-3 Conjecture, posed by Karoński, Łuczak, and Thomason [5], asserted that every nice graph is $\{1,2,3\}$-edge-weightable. After remaining open for two decades, this conjecture was finally confirmed by Keusch [6] in 2024. The property of being $\{1,2,3\}$-edge-weightable corresponds precisely to universal 3-edge-weightability restricted to the specific set $\{1,2,3\}$. For regular graphs, the 1-2-3 result can be extended via an affine transformation to show that they are $\{a,b,c\}$-edge-weightable for any arithmetic progression $\{a,b,c\}$ (where ba=cbb-a=c-b). However, this approach fails for general (non-arithmetic) three-element sets, leaving open the question of whether regular graphs are universally 3-edge-weightable.
@@
Various strengthenings of the 1-2-3 Conjecture have been investigated. A graph is edge-weight $k$-choosable if for any assignment $L$ that gives each edge $e$ a list $L(e)$ of $k$ real numbers, there exists a proper edge weighting $w$ with $w(e)\in L(e)$ for all $e\in E$. The list version of the 1-2-3 Conjecture [1] asks whether every nice graph is edge-weight $3$-choosable. A stronger concept is $(1,3)$-choosability [7]: a graph is $(1,3)$-choosable if, for an arbitrary vertex weighting $f:V\rightarrow\mathbb{R}$ and any assignment of $3$-element lists to edges, there exists an edge weighting $w$ using weights from the respective lists such that $f(u)+d_{w}(u)\neq f(v)+d_{w}(v)$ for every edge $uv\in E$.
These properties form a clear hierarchy:
$(1,3)\text{-choosable}\ \Rightarrow\ 3\text{-choosable}\ \Rightarrow\ \text{universally 3-edge-weightable}\ \Rightarrow\ \{1,2,3\}\text{-edge-weightable}.$
@@
Progress on the strongest property, $(1,3)$-choosability, has been steady. Cao [3] proved that every nice graph is $(1,17)$-choosable, improved by Zhu [8] to $(1,5)$-choosability. Consequently, every nice graph is edge-weight $5$-choosable. However, it remains open whether every nice graph is $(1,3)$-choosable, or even edge-weight $3$-choosable. Partial results show that several graph classes are $(1,3)$-choosable, including complete graphs and complete bipartite graphs [1], generalized Halin graphs [12], graphs with maximum average degree less than $\frac{11}{4}$ [9], dense Eulerian graphs [13], and Cartesian products of paths and cycles [11].
Between the weakest property (settled $\{1,2,3\}$-edge-weightability) and the stronger open problem of $3$-choosability lies the natural question of universal 3-edge-weightability:
Problem 1.1 (Universal 3-edge-weightability).
Is every nice graph universally 3-edge-weightable?
In other words, can the fixed set $\{1,2,3\}$ be replaced by an arbitrary 3-element set? Our result (Corollary 2) shows that regular graphs are universally 3-edge-weightable, while Problem 1 asks whether this holds for all nice graphs.
A related but different line of research concerns total weightings, where weights are assigned to both vertices and edges. For this total weighting problem, the 1-2 Conjecture [10] asks whether every nice graph is $\{1,2\}$-total-weightable. Notably, for regular graphs a stronger universal 2-total-weightability has been established [4], which serves as motivation for the present work on universal 3-edge-weightability.
Our contribution
In this paper, we demonstrate that the independent set decomposition framework in [4] can be adapted from total weighting to edge weighting, and can furthermore handle universal 3-edge-weighting. Our main technical result is the following:
@@
For any positive real numbers $d_{1},d_{2}$ with $0<d_{1}<d_{2}$, every $k$-regular graph with $k\geq 3$ is $\{-d_{1},0,d_{2}\}$-edge-weightable.
A simple scaling and shifting argument yields the following corollary from Theorem 1.
@@
For $k=2$, a $2$-regular graph consists of disjoint cycles. Graphs with maximum average degree less than $\frac{11}{4}$ are known to be $(1,3)$-choosable [9]. Taking the vertex weighting $f\equiv 0$ and giving every edge the same list $\{a,b,c\}$, $(1,3)$-choosability implies the existence of a proper $\{a,b,c\}$-edge weighting. Thus we may assume $k\geq 3$.
It suffices to show that $G$ admits a proper $\{-(b-a),0,c-b\}$-edge weighting $w^{\prime}$. Indeed, if we then define $w=w^{\prime}+b$ (adding $b$ to every edge weight), the regularity of $G$ ensures that $w$ is a proper $\{a,b,c\}$-edge weighting.
@@
The result of Corollary 2 places regular graphs strictly between the settled 1-2-3 Conjecture and the open $3$-choosability question in the hierarchy of edge-weighting properties.
Methodological remarks
Our proof of Theorem 1 adapts the independent-set decomposition framework from total-weightings to edge-weightings while addressing several distinctive challenges.
1. Absence of vertex-weight adjustments. In total weighting, vertex weights provide an independent degree of freedom that can be used to resolve conflicts. Edge weighting lacks this flexibility. To compensate, after partitioning the vertices into special independent sets $I_{0},I_{1},\ldots,I_{\ell}$, we employ the maximum independent set $I_{0}$ as a global regulator. We systematically weight $G[\cup_{i\geq 1}I_{i}]$ with $\{0,d_{2}\}$, and assign weight $d_{1}$ to selected edges through matchings between $I_{0}$ and $\cup_{i\geq 1}I_{i}$. This creates the separation $d_{w}(v)<d_{w}(u)$ for all $v\in I_{i}$ and $u\in I_{j}$ with $1\leq i<j$, and separation $d_{w}(x)<d_{w}(z)$ for all $v\in I_{0}$ and $z\in\cup_{i\geq 2}I_{i}$. The separation prevents conflicts between $I_{0}$ and $G[\cup_{i\geq 2}I_{i}]$, and conflicts in $G[\cup_{i\geq 1}I_{i}]$.
2. Different behavior under affine transformations. For total weighting on $k$-regular graphs, any proper $\{a_{0},b_{0}\}$-total weighting $w$ can be affinely transformed into a proper $\{a,b\}$-total weighting $w^{\prime}$. The transformation $w^{\prime}\mapsto\frac{b-a}{b_{0}-a_{0}}(w-a_{0})+a$, applied uniformly to vertex and edge weights, yields
$d_{w^{\prime}}(v)=\frac{b-a}{b_{0}-a_{0}}d_{w}(v)+(k+1)\left(a-\frac{b-a}{b_{0}-a_{0}}a_{0}\right),$
which preserves properness because the additive term is constant across vertices. Proving the $\{0,1\}$-case therefore suffices for the universal $\{a,b\}$-property.
For edge weighting, affine transformations are more restrictive. An affine map $f(x)=\alpha x+\beta$ preserves the ratio $\frac{f(b_{0})-f(a_{0})}{f(c_{0})-f(b_{0})}=\frac{b_{0}-a_{0}}{c_{0}-b_{0}}$. Thus $\{f(a_{0}),f(b_{0}),f(c_{0})\}$ must have the same spacing ratio as $\{a_{0},b_{0},c_{0}\}$. In particular, $\{1,2,3\}$ can only be mapped to other arithmetic progressions. This limitation means that a single fixed 3-element set cannot generate all 3-element sets by affine transformation.
3. Handling arbitrary 3-element sets directly. Keusch’s theorem guarantees a proper $\{1,2,3\}$-edge weighting for every nice graph. For regular graphs, we extend this result to arbitrary 3-element set $\{a,b,c\}$.
For arithmetic progressions $\{a,b,c\}$ (ba=cbb-a=c-b), scaling reduces the problem to $\{1,2,3\}$, which is covered by Keusch’s theorem.
For non-arithmetic 3-element set, we first apply an affine transformation to obtain an equivalent 3-element set of the form $\{-d_{1},0,d_{2}\}$ with $0<d_{1}<d_{2}$. Theorem 1 then constructs a proper $\{-d_{1},0,d_{2}\}$-edge weighting. The construction works uniformly for all $d_{1},d_{2}$ satisfying $0<d_{1}<d_{2}$, thereby covering all non-arithmetic 3-element sets.
4. Two-phase construction. Theorem 1 is proved via a two-phase procedure.
Phase 1 (Global separation). Using matchings between $I_{0}$ and higher layers, and among the higher layers themselves, we establish the weighted degree separation $d_{w}(x)<d_{w}(z)$ for all $v\in I_{0}$ and $z\in\cup_{i\geq 2}I_{i}$, and $d_{w}(v)<d_{w}(u)$ for all $v\in I_{i}$ and $u\in I_{j}$ with $1\leq i<j$. Remaining conflicts occur only in the bipartite subgraph $G[I_{0},I_{1}]$.
Phase 2 (Local conflict resolution). We classify vertices in $I_{1}$ that are incident to conflicts into three types based on their number of neighbors in $I_{0}$: $X_{3^{+}}$ (at least three neighbors), $X_{2}$ (exactly two neighbors), and $X_{1}$ (exactly one neighbor). Each type is processed with a specific weighting scheme that eliminates conflicts while preserving the separation achieved in Phase 1.
5. Consequences. The two-phase construction shows that regular graphs admit proper 3-edge-weightings for all sets of distinct real numbers, not only for arithmetic progressions. This extends the conclusion of the 1-2-3 Conjecture from the fixed set $\{1,2,3\}$ to arbitrary weight sets of size three, for the class of regular graphs.
Organization
Section 2 is devoted to the proof of Theorem 1. The proof is divided into two main phases: the initial weighting phase (Phase 1) and the conflict-resolution phase (Phase 2).
@@ 2. Proof of Theorem 1
We will use the following two lemmas.
@@
[2] Each $d$-degenerate graph is $(d+1)$-colorable.
@@
[4] Let $G=(V,E)$ be a graph, $I$ be a maximum independent set of $G$, and $I^{\prime}$ be an independent set of $G[V\setminus I]$. Then the bipartite graph $G[I,I^{\prime}]$ admits a matching that saturates $I^{\prime}$.
@@
Let $G=(I_{0}\cup I_{1},E)$ be a bipartite graph with minimum degree at least three, and $d_{1},d_{2}$ be two real numbers with $0<d_{1}<d_{2}$. Then $G$ admits a proper $\{-d_{1},0,d_{2}\}$-edge weighting.
@@
We would define an edge weighting such that each vertex in $I_{1}$ has weighted degree in $\{-3d_{1},d_{2}\}$ and each one in $I_{0}$ has weighted degree distinct from that in $\{-3d_{1},d_{2}\}$.
In the beginning, every edge is weighted with 0. Let $I_{1}=\{x_{i}:i=1,2,\ldots,p\}$.
First, label three of the edges incident to $x_{1}$ with $d_{1}$. Then $d_{w}(x_{1})=-3d_{1}$, and for each $y\in I_{0}$, $d_{w}(y)\in\{0,-d_{1}\}$; note that $\{0,-d_{1}\}$ is disjoint from $\{-3d_{1},d_{2}\}$.
@@
(ii) for each vertex $y$ in $I_{0}$, $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1},d_{2}\}$;
@@
Case 1: If $x_{i}$ has a neighbor $y^{\ast}$ with $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$, let $w(x_{i}y^{\ast})=d_{2}$. After this assignment, $d_{w}(x_{i})=d_{2}$, and the weighted degree of $y^{\ast}$ increases by $d_{2}$. Since before the assignment $d_{w}(y^{\ast})<-d_{1}$, after adding $d_{2}$ we have $d_{w}(y^{\ast})<d_{2}-d_{1}$. Moreover, because originally $d_{w}(y^{\ast})\neq-d_{2}-3d_{1}$, the new value satisfies $d_{w}(y^{\ast})\neq-3d_{1}$. Thus $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, and one can verify that conditions (i), (ii), (iii) are maintained for $i+1$.
Case 2: If each neighbor of $x_{i}$ has a weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$, let $y_{1},y_{2},y_{3}$ be three of them. Let $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$. Then $d_{w}(x_{i})=-3d_{1}$ and each neighbor of $x_{i}$ has a weighted degree in $2d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1},-d_{2}-4d_{1}\}$ (this set is disjoint with $\{-3d_{1},d_{2}\}$), and one can check that (i), (ii), (iii) hold for $i+1$.
Then finally, each vertex in $I_{0}$ has weighted degree in $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1},d_{2}\}$, while each one in $I_{1}$ has weighted degree in $\{-3d_{1},d_{2}\}$, and we obtain a proper $\{-d_{1},0,d_{2}\}$-edge weighting. ∎
@@
Partition. We partition $V$ into independent sets as follows. Choose a maximum independent set $I_{0}$ in $G$. For $i\geq 1$, let $I_{i}$ be a maximum independent set in $G_{i}$, where $G_{i}=G[V\setminus(\cup_{j<i}I_{j})]$. Let $\ell$ be the maximum integer such that $I_{\ell}\neq\emptyset$. Then $I_{\ell}$ is an independent set, and $1\leq\ell\leq k+1$ since $G$ is nonempty and each vertex in $I_{\ell}$ has a neighbor in $I_{i}$ for each $i<\ell$. If $\ell=1$, the conclusion holds from Lemma 3. So suppose $2\leq\ell\leq k+1$.
We will construct a proper $\{-d_{1},0,d_{2}\}$-edge weighting $w$ where construction proceeds in two phases. In the weighting process, for a vertex $v$ in $I_{i}$ ($i\in[1,\ell]$), we say $v$ is:
  • Type I if $d_{w}(v)=(i-1)d_{2}$, and $v$ is incident to exactly (i1)(i-1) edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$;
@@
  • hungry if $v\in I_{i}$ for some $i\geq 2$, $d_{w}(v)=jd_{2}$ for some $j\leq i-2$ and $v$ is incident to exactly $j$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$;
Goal. We will make sure $w$ satisfies the following conditions:
@@
    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=-3d_{1}$, $x$ has at least two neighbors in $I_{0}$, and every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\neq-3d_{1}$;
@@
    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=0$, every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\neq 0$;
Note that in (1), $\max\{(i-1)d_{2},id_{2}-2d_{1}\}<id_{2}-d_{1}$ because $d_{1}<d_{2}$. Moreover, for $1\leq i<j$, one has $id_{2}-d_{1}<\min\{(j-1)d_{2},jd_{2}-2d_{1}\}$. So (1), (2), (3) ensure there is no conflict in $G[\cup_{i\geq 1}I_{i}]$, and here is no conflict between $I_{0}$ and $\cup_{i\geq 2}I_{i}$. Clearly, (3), (4), (5), (6) ensure there is no conflict between $I_{0}$ and $I_{1}$.
The construction proceeds in two phases. Phase 1 ensures condition (1), (2),(3) hold; Phase 2 ensures conditions (4), (5), (6) hold, with (1), (2),(3) remaining valid.
@@
Step 1 By Lemma 2, there exists a matching $M_{\ell-1,\ell}$ in the bipartite graph $G[I_{\ell-1},I_{\ell}]$ that saturates $I_{l}$. Label each edge in $M_{\ell-1,\ell}$ with $d_{2}$. Then when $\ell=2$, each vertex in $I_{2}$ is Type I (has weighted degree $d_{2}=(2-1)d_{2}$); when $\ell\geq 3$, each vertex in $I_{\ell}$ is hungry (has weighted degree $d_{2}\leq(\ell-2)d_{2}$). Let $I^{\ast}_{\ell-1}$ be the set of vertices in $I_{\ell-1}$ that are saturated by $M_{\ell-1,\ell}$. Then the following holds.
@@
  • When $\ell=2$, by Lemma 2, there exists a matching $M^{\ast}_{0,\ell-1}$ in the bipartite graph $G[I_{0},I^{\ast}_{1}]$ that saturates $I^{\ast}_{1}$; label each edge in $M^{\ast}_{0,1}$ with $d_{1}$, which makes each vertex in $I^{\ast}_{1}$ Type II (has weighted degree $d_{2}-d_{1}$), and each one in $I_{1}\setminus I^{\ast}_{1}$ Type I (has weighted degree $0=(1-1)d_{2}$).
Then when $\ell=2$, one can check that (1’), (2’) already hold.
Step 2 When $\ell\geq 3$, we process the indices $i=\ell-2,\ell-3,\ldots,1$ in decreasing order. At the start of step $i$, we assume:
@@
    every vertex in $I_{\alpha}$ for each $\alpha\geq i+1$ is Type I, Type II or hungry;
@@
    for $\alpha\geq i+2$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta\in[i+1,\alpha-1]$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.

### v2

# Regular graphs are universally 3-edge-weightable

@@ Abstract.
A graph is universally $k$-edge-weightable if for every $k$-element set $Q\subset\mathbb{R}$, it admits a proper $Q$-edge weighting. The settled 1-2-3 conjecture implies that for any arithmetic progression $\{a,b,c\}$, every nice regular graph has a proper $\{a,b,c\}$-edge weighting. We prove that this remains valid for all 3-element set $\{a,b,c\}$ with $c-b\neq b-a$. Consequently, every nice regular graph is universally $3$-edge-weightable.
@@ 1. Introduction
All graphs considered are simple, finite, and undirected. For notation or terminology not defined here, we refer to [1]. A graph is nice if it contains no $K_{2}$ component. For $G=(V,E)$ and $Q\subset\mathbb{R}$, a $Q$-edge weighting is a mapping $w:E\to Q$. When $Q=\{1,2,\ldots,k\}$, it is called a $k$-edge weighting. The weighted degree of $v$ is $d_{w}(v)=\sum_{e\ni v}w(e)$. An edge $uv$ is a conflict if $d_{w}(u)=d_{w}(v)$; $w$ is proper if it has no conflicts.
The 1-2-3 conjecture, posed by Karoński, Łuczak, and Thomason [2], asserts that every nice graph has a proper $\{1,2,3\}$-edge weighting. This conjecture is a local version of the classical–and still active–irregularity strength problem [3, 4, 6, 7, 5, 8, 9]. Research on the 1-2-3 conjecture has progressed along several fronts. Upper bounds for general graphs were successively lowered: from 30 [10], to 16 [11], 13 [12], 5 [13], and finally to 3 [14] after an intermediate improvement to 4 [15]. For specific classes, it was verified for dense graphs [16], $d$-regular graphs with large $d$ [17], graphs with $\delta=\Omega(\log\Delta)$ [18], and shown to hold asymptotically almost surely for random graphs [11]. Complexity results include NP-completeness for deciding proper $\{1,2\}$-edge weightings [19] and a polynomial-time algorithm for bipartite graphs [20]. The conjecture was affirmatively resolved by Keusch [14], leading to the 1-2-3 theorem. For more 1-2-3 conjecture related problems, refer to [21, 22, 23, 24, 25, 26, 27] and the references cited there.
@@
Various strengthenings of the 1-2-3 conjecture have been studied. The list version [28] asks whether every nice graph is edge-weight $3$-choosable, i.e. for any assignment of $3$-element lists $L(e)$ to edges, there exists a proper edge weighting $w$ with $w(e)\in L(e)$ for all $e$. Wong and Zhu [29] conjectured a stronger property called $(1,3)$-choosability, where each vertex receives a fixed weight (from a $1$-element list) and each edge receives a weight from a $3$-element list; the weighted degree of a vertex is then $d_{w}(v)=w(v)+\sum_{e\ni v}w(e)$. These properties form a clear hierarchy:
$(1,3)\text{-choosable}\Rightarrow\text{edge-weight 3-choosable}\Rightarrow\text{the 1-2-3 theorem}.$
@@
Progress on the strongest property, $(1,3)$-choosability, has been steady: Cao [30] proved that every nice graph is $(1,17)$-choosable, improved by Zhu [31] to $(1,5)$-choosability. Consequently, every nice graph is edge-weight $5$-choosable. However, it remains open whether every nice graph is $(1,3)$-choosable, or even edge-weight $3$-choosable. Partial results show that several graph classes are $(1,3)$-choosable, including complete graphs and complete bipartite graphs [28], generalized Halin graphs [32], graphs with maximum average degree less than $\frac{11}{4}$ [33], dense Eulerian graphs [34], and Cartesian products of paths and cycles [35].
The 1-2-3 theorem guarantees a proper $\{1,2,3\}$-edge weighting for every nice graph. A natural weakening of the unsolved $(1,3)$-choosability conjecture and edge-weight $3$-choosability conjecture, is to ask whether this extends to any three-element set $Q\subset\mathbb{R}$, not just $\{1,2,3\}$. We call a graph universally $3$-edge-weightable if it admits a proper $Q$-edge weighting for every such $Q$.
For regular graphs, affine transformation from the 1-2-3 theorem immediately gives a proper $\{a,b,c\}$-edge weighting for any arithmetic progression $\{a,b,c\}$ (where ba=cbb-a=c-b). The non-arithmetic case, however, requires genuinely new constructions. This paper provides the first such construction, which implies that every regular graph is universally $3$-edge-weightable.
@@
For any positive real numbers $d_{1},d_{2}$ with $0<d_{1}<d_{2}$, every $k$-regular graph with $k\geq 3$ admits a proper $\{-d_{1},0,d_{2}\}$.
@@
For $k=2$, a $2$-regular graph consists of disjoint cycles. Graphs with maximum average degree less than $\frac{11}{4}$ are known to be $(1,3)$-choosable [33], implying the existence of a proper $\{a,b,c\}$-edge weighting. Thus we may assume $k\geq 3$.
It suffices to show that $G$ admits a proper $\{-(b-a),0,c-b\}$-edge weighting $w^{\prime}$. Indeed, if we define $w=w^{\prime}+b$ (that is, we add $b$ to every edge weight), the regularity of $G$ ensures that $w$ is a proper $\{a,b,c\}$-edge weighting.
@@
This settles the regular case completely. Whether universal $3$-edge-weightability holds for all nice graphs remains open:
Problem 1.1 (Universal 3-edge-weightability).
Is every nice graph universally $3$-edge-weightable?
Theorem 1 is proved in Section 2.
@@ 2. Proof of Theorem 1
We will use the following three lemmas.
@@
[1] Each $i$-degenerate graph $(i\geq 1)$ is $(i+1)$-colorable.
@@
[36] Let $G=(V,E)$ be a graph, $I_{0}$ be a maximum independent set of $G$, and $I_{1}$ be an independent set of $G[V\setminus I_{0}]$. Then the bipartite graph $G[I_{0},I_{1}]$ admits a matching saturating $I_{1}$.
@@
Let $G=(I_{0}\cup I_{1},E)$ be a bipartite graph with $d(v)\geq 3$ for each $v\in I_{1}$; let $d_{1}$ and $d_{2}$ be two real numbers with $0<d_{1}<d_{2}$. Then $G$ admits a proper $\{-d_{1},0,d_{2}\}$-edge weighting.
@@
We construct an edge weighting $w$ such that every vertex in $I_{1}$ attains a weighted degree in $\{-3d_{1},d_{2}\}$ and each one in $I_{0}$ has weighted degree in $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
Initially, we assign weight 0 to every edge. Let $I_{1}=\{x_{i}:i=1,2,\ldots,p\}$.
First, label three of the edges incident to $x_{1}$ with $d_{1}$. Then $d_{w}(x_{1})=-3d_{1}$, and for each $y\in I_{0}$, $d_{w}(y)\in\{0,-d_{1}\}\subseteq(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
@@
(ii) for each vertex $y$ in $I_{0}$, $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$;
@@
Case 1: If $x_{i}$ has a neighbor $y^{\ast}$ with $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$, let $w(x_{i}y^{\ast})=d_{2}$. After this assignment, $d_{w}(x_{i})=d_{2}$, $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, and each $y\in I_{0}\setminus\{y^{\ast}\}$ still satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. Thus, conditions (i)-(iii) are maintained for $i+1$.
Case 2: If each neighbor of $x_{i}$ has a weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$, let $y_{1},y_{2}$ and $y_{3}$ be three of them. Set $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$. Then $d_{w}(x_{i})=-3d_{1}$. For each $j=1,2,3$, $d_{w}(y_{j})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$, which is a subset of $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. What is more, each $y\in I_{0}\setminus\{y_{1},y_{2},y_{3}\}$ still satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. Thus, conditions (i)-(iii) hold for $i+1$.
Consequently, each vertex in $I_{0}$ attains a weighted degree in $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, while each one in $I_{1}$ has weighted degree in $\{-3d_{1},d_{2}\}$, and we obtain a proper $\{-d_{1},0,d_{2}\}$-edge weighting. ∎
@@
Partition. We partition $V$ into independent sets as follows. Choose a maximum independent set $I_{0}$ in $G$. For $i\geq 1$, let $I_{i}$ be a maximum independent set in $G_{i}$, where $G_{i}\triangleq G[V\setminus(\cup_{j<i}I_{j})]$. Let $\ell$ be the maximum integer such that $I_{\ell}\neq\emptyset$. Then $I_{\ell}$ is an independent set, and $1\leq\ell\leq k+1$ since $G$ is nonempty and each vertex in $I_{\ell}$ has a neighbor in $I_{i}$ for each $i<\ell$. If $\ell=1$, the conclusion holds from Lemma 3. So suppose $2\leq\ell\leq k+1$.
We will construct a proper $\{-d_{1},0,d_{2}\}$-edge weighting $w$ in two phases. In the weighting process, for $i=1,2,\ldots,\ell,$ a vertex $v$ in $I_{i}$ is called:
  • Type I if $d_{w}(v)=(i-1)d_{2}$, and $v$ is incident to exactly i1i-1 edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$;
@@
  • hungry if $v\in I_{i}$ for some $i\geq 2$, $d_{w}(v)=pd_{2}$ for some $p\leq i-2$ and $v$ is incident to exactly $p$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$.
Goal. We will make sure $w$ finally satisfies the following conditions:
@@
    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=-3d_{1}$, $x$ has at least two neighbors in $I_{0}$, and every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$;
@@
    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=0$, every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{0\}$.
Note that in (1), $\max\{(i-1)d_{2},id_{2}-2d_{1}\}<id_{2}-d_{1}$ because $d_{1}<d_{2}$. Moreover, for $1\leq i<j$, one has $id_{2}-d_{1}<\min\{(j-1)d_{2},jd_{2}-2d_{1}\}$. So (1)-(3) ensure there is no conflict in $G[\cup_{i\geq 1}I_{i}]$, and there is no conflict between $I_{0}$ and $\cup_{i\geq 2}I_{i}$. Clearly, (3)-(6) ensure there is no conflict between $I_{0}$ and $I_{1}$.
The construction proceeds in two phases. Phase 1 ensures condition (1)-(3) hold; Phase 2 ensures conditions (4)-(6) hold, with (1)-(3) remaining valid.
@@
Clearly, the validity of conditions (1’)-(3’) guarantees that (1)-(3) hold.
Step 1 By Lemma 2, there exists a matching $M_{\ell-1,\ell}$ in the bipartite graph $G[I_{\ell-1},I_{\ell}]$ that saturates $I_{l}$. Weight each edge in $M_{\ell-1,\ell}$ with $d_{2}$. Then when $\ell=2$, each vertex in $I_{2}$ is Type I (has weighted degree $d_{2}=(2-1)d_{2}$); when $\ell\geq 3$, each vertex in $I_{\ell}$ is hungry (has weighted degree $d_{2}\leq(\ell-2)d_{2}$). Let $I^{\ast}_{\ell-1}$ be the set of vertices in $I_{\ell-1}$ that are saturated by $M_{\ell-1,\ell}$.
  • When $\ell=2$, by Lemma 2, there exists a matching $M^{\ast}_{0,1}$ in the bipartite graph $G[I_{0},I^{\ast}_{1}]$ that saturates $I^{\ast}_{1}$; weight each edge in $M^{\ast}_{0,1}$ with $d_{1}$, which makes each vertex in $I^{\ast}_{1}$ Type II (has weighted degree $d_{2}-d_{1}$), and each one in $I_{1}\setminus I^{\ast}_{1}$ Type I (has weighted degree $0=(1-1)d_{2}$). Then (1’) and (2’) already hold in this case.
@@
Step 2 When $\ell\geq 3$, we process the indices $i=\ell-2,\ell-3,\ldots,1$ in decreasing order. Right before proceeding $i$, we assume:
@@
    for each $\alpha\geq i+1$, every vertex in $I_{\alpha}$ is Type I, Type II or hungry;
@@
    for $\alpha\geq i+2$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta=i+1,\ldots,\alpha-1$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.